MDM CCD Noise and OWL Run Failure

Thursday, February 7, 2013 3:21 PM

Problem(s) Encountered:

Still have variable noise in bias.

OWL lost connection to controller, so we power cycled and restored..

This happened after we used `ABORT''- coincidence? the lost

connection the previous night might have been after an ABORT as

well but we don''t remember.


In an attempt to mitigate the variable electronic noise yesterday, Tony and I swapped the electronics box from the gold box over to the silver box that I have normally used with these detectors.  This was done as there is possibly some confusion over whether or not the gold box is suitable for detectors other than Wilbur (the silver box had previously been in use with Nellie at the 2.4m).  Ran a couple bias frames after this switch and all looked well.  Fast forward to last night''s report:

Ran numerous bias frames this morning.  I then powered down the CCD electronics and switched the power away from the filter box, connecting it instead by the mirror cover electronics.  I then took more bias frames.  I don''t see any discernible differences between using the filter and not using it.  The first frame after powering up the detector appears to be about 5 counts on average lower than subsequent frames.  Not sure what causes this, just an observation.  

While there does appear to be some electronic noise intermittently in these frames, showing up as horizontal swaths of varying size, the effect is subtle.  Deltas in counts is very small, typically ~2-4 counts from what I can gather.  Not sure if this is typically considered to be "lost in the noise".

As for the OWL failure, it appears that the observers were able to get the program back up and running although I am unsure as to whether or not they had to power cycle the electronics.  Running frames today, I had no issue with OWL failing.